1. Three binary compounds composed of bromine and oxygen were decomposed and the mass of each element in a sample of each compound determined. The results were:

              Cpd 1: 7.21 g O and 72.0 g Br

              Cpd 2: 4.68 g O and 11.7 g Br

              Cpd 3: 5.43 g O and 18.1 g Br

a) Calculate the ratio of g Br/g O for each compound.

b) Are the ratios from (a) consistent with the Law of Multiple Proportions? Explain your answer using BOTH math and words.

c) If Cpd 2 has the formula BrO2 what are the formulas of the other two compounds?

2. Bromine atoms (Br) have a mass that is 4.21 times the mass of fluorine (F) atoms. A compound composed of the elements bromine and fluorine is found to have a mass ratio of 1.40 g Br/ g F.

  1. What is the smallest whole number ratio of the number of atoms of Br and F in the compound?

b) Is your answer to part (a) the only possible answer given the data  provided in the question? Explain fully.

3. The mineral galena is composed of the elements lead (Pb) and sulfur (S) and contains 6.45 g Pb/g S.

a) What is the % by mass of Pb in galena?

b) Another compound of lead and sulfur is found to contain 3.23 g Pb/g S. The chemical formula of galena is PbS. What is the formula of the other compound?
Answer #1

1. a) Cpd 1: 7.21 g O and 72.0 g Br

Ratio of g Br/g O = 72.0/7.21 = 9.9

Cpd 2: 4.68 g O and 11.7 g Br

Ratio of g Br/g O = 11.7/4.68 = 2.5

Cpd 3: 5.43 g O and 18.1 g Br

Ratio of g Br/g O = 18.1/5.43 = 3.33

b)

The law of multiple proportions states that :

If two elements form more than one compound between them, then the ratios of the masses of the second element which combine with a fixed mass of the first element will always be ratios of small whole numbers.

So let's check the mass of Bromine in the three compounds reacting with a fixed mass of Oxygen.

Cpd 1: 7.21 g O reacts with 72.0 g Br

Cpd 2: 4.68 g O reacts with 11.7 g Br 7.21 g O reacts with 18 g

Cpd 3: 5.43 g O reacts with 18.1 g Br 7.21 g O reacts with 24 g

Thus we can see, the mass of Br reacting with a fixed mass of O is 72 : 18 : 24 i.e. 12:3:4 which is consistent with the law of multiple proportions.

c)

Cpd 1: 7.21 g O and 72.0 g Br

number of moles of O = 7.21/16 = 0.45

number of moles of Br = 72/80 = 0.9

The ratio of moles of Br : O = 0.9 : 0.45 = 2 : 1.

The compound is Br2O

Cpd 3: 5.43 g O and 18.1 g Br

number of moles of O = 5.43/16 = 0.34

number of moles of Br = 72/80 = 0.23

The ratio of moles of Br : O = 0.34 : 0.23 = 2 : 3.

The compound is Br2O3